OpenStudy (anonymous):

find all the zeros of f(x)=x^4+x^3-23x^2-3x +60,if one of its zeors is root 3

OpenStudy (anonymous):

\[-\sqrt{3}\]

OpenStudy (anonymous):

when x is -5, 4, -1.732051, 1.7320508 actually it's two zeros with negative 3

OpenStudy (anonymous):

4,-5

OpenStudy (anonymous):

give me its solution

OpenStudy (anonymous):

may be you should do it in trail and error method. i did like that.

OpenStudy (cruffo):

If \[x = \sqrt{3}\] is a zero then it's conjugate is also a zero. So, \[ x = - \sqrt{3}\] is another zero. Multiply the factors, \[(x-\sqrt{3})(x+\sqrt{3}) = x^2-3\] Using long division of polynomials will lead to \[(x^2-3)(x^2+x-20) = (x^2-3)(x+5)(x-4)\] Giving the last two zeros of -5 and 4.